Lambda Computer Products competed for and won a contract to produce two prototype units of a new type of computer that is based on laser optics rather than on electronic binary bits. The first unit produced by Lambda took 5,000 hours to produce and required $250,000 worth of material, equipment usage, and supplies. The second unit took 3,500 hours and used $200,000 worth of material, equipment usage, and supplies. Labor is $30 per hour.

(a) Lambda was asked to present a bid for 10 additional units as soon as the second unit was completed. Production would start immediately. What would this bid be?

(b) Suppose that Lambda was asked to present a bid for an additional 10 units after 3 months delay since the second unit was completed. During this time, personnel and equipment were reassigned to other projects. What would be the total production costs for the additional 10 units? (--> When you consider the forgetting effects for this problem, make sure that the PRODUCTION UNIT must be an integer, i.e., if you get any non-integer values from the equation, it needs to be rounded to the nearest integer value.)
Answer #1

Answer a=

LR = 3500/5000 = 70%

cumulative factor frelated to labor for 12 units= 5.501

cumulative factor frelated to labor for 2 units= 1.700

cumulative factor = 5.501 – 1.700 = 3.801

Labor cost for 10 more units = 5,000(3.801)(30) = $570,150

For Material:

LR = 200000/250000 = 80%

cumulative factor in case of material for unt 12 = 7.227

cumulative factor in case of material for unt 2 = 1.800

Cumulative factor = 7.227 – 1.800 = 5.427

Total cost fo material for aditioanl 10 units = 250,000(5.427) = $1,356,750

Total Cost - total labor cost+ total materail cost

Total cost = $570,150 + $1,356,750 = $1,926,900

So the bid will be for $1,926,900

Answer b The minimum bid will be as determined in part a that is $1,926,900

If we assume that no benefits are assumed to be gained from unit 1 and unit 2

So the labor cost =

Cumulative factor=4.931

Labor cost = 4.931*(5,000)*(30)  =$739,650

So the material cost =

Cumulative factor=6.315

Materail cost =  6.315*($250,000) =$1,578,750

Total cost=$739,650+$1,578,750

Total cost= $2,318,400

So the minimum bid will be $1,926,900 and the maximum bid can be $2,318,400